OpenStudy (anonymous):

Need help!!! please!! The present value of the building in the downtown area is given by the function : P(t) = 300,000e^-0.09t+square root of t, over 2. For t = and grater than 0 and t = and less than 10. Find the optimal present value of the building. ( Hint : the slope of the tangent line is 0.)

OpenStudy (anonymous):

I am very confusing

OpenStudy (anonymous):

e^-0.09(o)+square root of 0 over 2 = e^0 = 1 right ?????

OpenStudy (anonymous):

P(0)= 300,000 (1) = 300,000 dollars ???

OpenStudy (anonymous):

I hope somebody can help me tomorrow!!

OpenStudy (anonymous):

Derive the present Value function to get P'(t). Make the derivative = zero and solve for t. Take the solution within the interval o<t<10. Just substitute the value of that t back into Present value function. I think that's what you do.

OpenStudy (anonymous):

Thank you mebs !!!!

OpenStudy (anonymous):

yw